OpenStudy (anonymous):

f(x) = 2x2 - 12x + 2 [0, 8] find the absolute max and absolute min of f on the given interval

OpenStudy (anonymous):

Take the derivative.=> 4x -12 set it equal to 0 to find critical points. 0 = 4x -12 => x = 3

OpenStudy (anonymous):

f'(x)=4x-12---> x= 3 f(0)=2 f(8)=2(64)-12(8)+2=....

OpenStudy (anonymous):

Try x=3, and our boundary points 0 , and 8 into the original equation. and figure out which gives the biggest value and which gives the smallest value.